### v1

@@ 1 Introduction
However, a fundamental limitation of the single lattice approach was revealed by [3], which proved that any algorithm based on a single rank-1 lattice fails to achieve the optimal $L_{2}$ error rate in periodic Sobolev spaces, including the Korobov space as a special case. While the optimal rate for these spaces should be of order $N^{-\alpha}$ up to logarithmic factors, any deterministic single lattice approach can attain the rate at most of order $N^{-\alpha/2}$. We point out that this limitation is, in fact, inherently related to classical results in the theory of number-theoretic methods; similar lower bound arguments can be traced back to the work of Korobov [16] and the monograph by Hua and Wang [10, Theorems 9.14 and 9.15]. For the reader’s convenience, we summarize these classical lower bound arguments in Appendix B.
@@
    We establish that both the deterministic and randomized error bounds are strongly tractable under suitable summability conditions on the general weights. In the case of product weights, the summability condition we obtain is weaker than that required in [8, Theorem 15.11] when $\alpha<2$.
@@ 3.2 Multiple lattice-based algorithms
The total computational complexity for determining all $U_{t}(\boldsymbol{k})$ for $t\in\{1,\ldots,L\}$ and $\boldsymbol{k}\in\mathcal{A}_{d}$ is of $\mathcal{O}(dL|\mathcal{A}_{d}|\log(L|\mathcal{A}_{d}|))$. This efficiency ensures that the precomputation of the aliasing-free index sets remains feasible even in high dimensions.
@@
$A_{N}^{\mathrm{mult}}(f)(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\widehat{f}^{\mathrm{mult}}_{N}(\boldsymbol{k})\,e^{2\pi i\boldsymbol{k}\cdot\boldsymbol{x}}.$ (5)
@@ 4.1 The worst-case $L_{\infty}$ error
with probability at least $1-\delta$. Here, the constant $C_{c,\delta,\beta,\lambda}$ depends only on $c,\delta,\beta,\lambda$ and is independent of $N,\alpha$, and the weights $\boldsymbol{\gamma}$. In particular, as $\lambda\to 1/\alpha$ and $\beta\to 0$, the convergence rate approaches arbitrarily close to $\mathcal{O}(N^{-\alpha+1/2})$.
@@ 4.2 The worst-case root mean squared $L_{2}$ error
$\mathrm{err}_{2}^{\text{rand}}(A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}},H_{d,\alpha,\boldsymbol{\gamma}})\leq\frac{\sqrt{L^{2}+1}}{M},$
@@
$\int_{[0,1]^{d}}\|f-A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(f)\|_{L_{2}}^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq\|f\|_{d,\alpha,\boldsymbol{\gamma}}^{2}\frac{L^{2}+1}{M^{2}}$
@@
$\int_{[0,1]^{d}}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\left|\widehat{f}(\boldsymbol{k})-\widehat{f}_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(\boldsymbol{k})\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq L^{2}\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}.$
@@
where the third equality follows from the character property of the lattice (Lemma 2.3), and the last equality follows from the change of variables $\boldsymbol{\ell}=\boldsymbol{h}-\boldsymbol{k}$. By the Cauchy–Schwarz inequality and the fact that $U^{2}_{t}(\boldsymbol{k})=U_{t}(\boldsymbol{k})$, we have
@@
$\leq\int_{[0,1]^{d}}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{L}{(\xi(\boldsymbol{k}))^{2}}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\left|\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})\,e^{2\pi i\boldsymbol{\ell}\cdot\boldsymbol{\Delta}}\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}$
$=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{L}{(\xi(\boldsymbol{k}))^{2}}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\int_{[0,1]^{d}}\left|\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})\,e^{2\pi i\boldsymbol{\ell}\cdot\boldsymbol{\Delta}}\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}$
$=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{L}{(\xi(\boldsymbol{k}))^{2}}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}$
$\leq L\sum_{t=1}^{L}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}U_{t}(\boldsymbol{k})\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}$
$=L\sum_{t=1}^{L}\sum_{\boldsymbol{k}\in\mathcal{A}_{d,t}}\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}.$
@@
Summing over $t$, we obtain the bound $L^{2}\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}$ as claimed.
@@
$\leq(L^{2}+1)\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}$
$\leq(L^{2}+1)\left(\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})|\widehat{f}(\boldsymbol{k})|^{2}\right)\sup_{\boldsymbol{k}\notin\mathcal{A}_{d}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})}$
$\leq\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}\frac{L^{2}+1}{M^{2}},$
@@
with probability at least $1-\delta$. Here, the constant $C^{\prime}_{c,\delta,\beta,\lambda}$ depends only on $c,\delta,\beta,\lambda$ and is independent of $N,\alpha$, and the weights $\boldsymbol{\gamma}$. In particular, as $\lambda\to 1/\alpha$ and $\beta\to 0$, the convergence rate approaches arbitrarily close to $\mathcal{O}(N^{-\alpha})$.
@@
Here, we considered the worst-case root mean squared $L_{2}$ error by incorporating a random shift. It is natural to ask whether similar optimal convergence rates and good tractability results, the latter of which shall be shown in the next subsection, can be achieved for the deterministic $L_{2}$ error without such a shift. As shown in the proof of Theorem 4.6, for an individual function $f\in H_{d,\alpha,\boldsymbol{\gamma}}$, the squared $L_{2}$ error for the randomized algorithm satisfies
$\int_{[0,1]^{d}}\|f-A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(f)\|_{L_{2}}^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}+L\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\sum_{t=1}^{L}U_{t}(\boldsymbol{k})\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2},$
@@ 4.3 Tractability
Finally, we discuss the tractability of the multiple lattice-based algorithms. We recall that the error bounds in Corollaries 4.4 and 4.7 hold under the assumption that
@@
$\geq 2\max_{j=1,\ldots,d}\max_{j\in u\subseteq[1{:}d]}\lfloor(\gamma_{u}M)^{1/\alpha}\rfloor\geq N_{\mathcal{A}_{d}},$
@@
Next, to show when the assumption $\eta\geq 4L_{\max}\log L_{\max}$ holds, let us consider the case $\delta=1/2$. From the definition of $L_{\max}$, by applying the inequality $\log(x+1)\leq x^{\beta}/\beta$, which holds for any $x\geq 0$ and $\beta\in(0,1)$ (here we take $\beta=1/3$), we have
@@
$\leq 36\left(\frac{c}{c-1}\right)^{8/3}\left(1+\frac{1}{\log 2}\right)^{4/3}\left(|\mathcal{A}_{d}|-1\right)^{4/9}.$
@@
$c\geq 36\left(\frac{c}{c-1}\right)^{8/3}\left(1+\frac{1}{\log 2}\right)^{4/3}.$
Numerically, this inequality is satisfied for $c>121.078$. It is important to note that the assumptions required for the error bounds in Corollaries 4.4 and 4.7 can be satisfied with $\delta=1/2$ and a constant $c>121.078$ independently of the dimension $d$, smoothness $\alpha$, and the weight parameters $\boldsymbol{\gamma}$.
We now state the conditions for strong polynomial tractability. Here, a multivariate problem is said to be strongly polynomially tractable if the number of function evaluations $N$ required to reduce the worst-case error less than or equal to $\varepsilon$ is independent of $d$, and is bounded by a polynomial of $\varepsilon^{-1}$. The following result is clear from Corollaries 4.4 and 4.7.
@@
When the weights $\boldsymbol{\gamma}$ have some specific structures, in particular, when they are either the product weights or the POD weights, the condition given in the above theorem can be simplified as follows. We refer to a recent work [25] for more refined analyses for the POD and SPOD weights cases.
@@
This completes the proof.
@@
$\sum_{|u|=k}\prod_{j\in u}x_{j}\leq\frac{1}{k!}\left(\sum_{j=1}^{\infty}x_{j}\right)^{k},$
@@
In [8, Theorem 15.11], it is claimed that, for product weights, strong polynomial tractability holds if there exists $\tau>1$ such that

### v2

@@ 1 Introduction
However, a fundamental limitation of the single lattice approach was revealed by [3], which proved that any algorithm based on a single rank-1 lattice fails to achieve the optimal $L_{2}$ error rate in periodic Sobolev spaces, including the Korobov space as a special case. While the optimal rate for these spaces should be of order $N^{-\alpha}$ up to logarithmic factors, any deterministic single lattice approach can attain a rate which is at most of order $N^{-\alpha/2}$. We point out that this limitation is, in fact, inherently related to classical results in the theory of number-theoretic methods; similar lower bound arguments can be traced back to the works of Smolyak [27] and Korobov [16], as well as the monograph by Hua and Wang [10, Theorems 9.14 and 9.15]. For the reader’s convenience, we summarize these classical lower bound arguments in Appendix B.
@@
    We establish that both the deterministic and randomized error bounds are independent of the dimension $d$ under suitable summability conditions on the general weights. In the case of product weights, the summability condition we obtain is weaker than that required in [8, Theorem 15.11] when $\alpha<2$.
@@ 3.2 Multiple lattice-based algorithms
To evaluate $U_{t}(\boldsymbol{k})$ efficiently for all $t\in\{1,\ldots,L\}$ and $\boldsymbol{k}\in\mathcal{A}_{d}$, we can employ a simple counting (or binning) strategy. For a fixed $t$, note that the values $\nu_{\boldsymbol{k},t}\coloneqq\boldsymbol{k}\cdot\boldsymbol{g}_{t}\pmod{n_{t}}$ are integers in $\{0,1,\ldots,n_{t}-1\}$. We first allocate an integer array of length $n_{t}$ initialized to all zeros. For each $\boldsymbol{k}\in\mathcal{A}_{d}$, we compute $\nu_{\boldsymbol{k},t}$ (which requires $\mathcal{O}(d)$ operations per frequency) and increment the corresponding entry in the array. This step takes $\mathcal{O}(d|\mathcal{A}_{d}|)$ operations. Then, a second pass over $\mathcal{A}_{d}$ allows us to determine $U_{t}(\boldsymbol{k})$ in $\mathcal{O}(|\mathcal{A}_{d}|)$ time: $U_{t}(\boldsymbol{k})=1$ if the array entry at index $\nu_{\boldsymbol{k},t}$ is exactly $1$, and 0 otherwise. Thus, the cost for a single lattice is $\mathcal{O}(d|\mathcal{A}_{d}|+n_{t})$. Summing this over all $L$ lattices, the total computational complexity becomes
$\mathcal{O}\left(dL|\mathcal{A}_{d}|+\sum_{t=1}^{L}n_{t}\right)=\mathcal{O}(dL|\mathcal{A}_{d}|+N).$
Since $N$ is bounded by $2cL_{\max}(|\mathcal{A}_{d}|-1)$ according to Proposition A.1 and $L_{\max}$ is chosen as in Algorithm 1, the overall complexity simplifies to $\mathcal{O}(dL_{\max}|\mathcal{A}_{d}|)=\mathcal{O}(d|\mathcal{A}_{d}|\log|\mathcal{A}_{d}|)$. This efficiency ensures that the precomputation of the aliasing-free index sets remains feasible even in high dimensions.
@@
$A_{N}^{\mathrm{mult}}(f)(\boldsymbol{x})=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\widehat{f}^{\mathrm{mult}}_{N}(\boldsymbol{k})\,e^{2\pi i\boldsymbol{k}\cdot\boldsymbol{x}},$ (5)
@@ 4.1 The worst-case $L_{\infty}$ error
with probability at least $1-\delta$. Here, the constant $C_{c,\delta,\beta,\lambda}$ depends only on $c,\delta,\beta,\lambda$ and is independent of $N,\alpha$, and the weights $\boldsymbol{\gamma}$. In particular, as $\lambda\to 1/\alpha$ and $\beta\to 0$, the convergence rate gets arbitrarily close to $\mathcal{O}(N^{-\alpha+1/2})$.
@@ 4.2 The worst-case root mean squared $L_{2}$ error
$\mathrm{err}_{2}^{\text{rand}}(A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}},H_{d,\alpha,\boldsymbol{\gamma}})\leq\frac{\sqrt{L+1}}{M},$
@@
$\int_{[0,1]^{d}}\|f-A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(f)\|_{L_{2}}^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq\|f\|_{d,\alpha,\boldsymbol{\gamma}}^{2}\frac{L+1}{M^{2}}$
@@
$\int_{[0,1]^{d}}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\left|\widehat{f}(\boldsymbol{k})-\widehat{f}_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(\boldsymbol{k})\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq L\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}.$
@@
where the third equality follows from the character property of the lattice (Lemma 2.3), and the last equality follows from the change of variables $\boldsymbol{\ell}=\boldsymbol{h}-\boldsymbol{k}$. By the Cauchy–Schwarz inequality and the fact that $U^{2}_{t}(\boldsymbol{k})=U_{t}(\boldsymbol{k})$, noting that the sum over $t$ contains exactly $\xi(\boldsymbol{k})$ non-zero terms for each $\boldsymbol{k}\in\mathcal{A}_{d}$, we have
@@
$\leq\int_{[0,1]^{d}}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{\xi(\boldsymbol{k})}{(\xi(\boldsymbol{k}))^{2}}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\left|\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})\,e^{2\pi i\boldsymbol{\ell}\cdot\boldsymbol{\Delta}}\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}$
$=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{1}{\xi(\boldsymbol{k})}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\int_{[0,1]^{d}}\left|\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})\,e^{2\pi i\boldsymbol{\ell}\cdot\boldsymbol{\Delta}}\right|^{2}\,\mathrm{d}\boldsymbol{\Delta}$
$=\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\frac{1}{\xi(\boldsymbol{k})}\sum_{t=1}^{L}(U_{t}(\boldsymbol{k}))^{2}\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}$
$\leq\sum_{t=1}^{L}\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}U_{t}(\boldsymbol{k})\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}$
$=\sum_{t=1}^{L}\sum_{\boldsymbol{k}\in\mathcal{A}_{d,t}}\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2}.$
@@
Summing over $t$, we obtain the bound $L\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}$ as claimed.
@@
$\leq(L+1)\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}$
$\leq(L+1)\left(\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})|\widehat{f}(\boldsymbol{k})|^{2}\right)\sup_{\boldsymbol{k}\notin\mathcal{A}_{d}}\frac{1}{r^{2}_{\alpha,\boldsymbol{\gamma}}(\boldsymbol{k})}$
$\leq\|f\|^{2}_{d,\alpha,\boldsymbol{\gamma}}\frac{L+1}{M^{2}},$
@@
with probability at least $1-\delta$. Here, the constant $C^{\prime}_{c,\delta,\beta,\lambda}$ depends only on $c,\delta,\beta,\lambda$ and is independent of $N,\alpha$, and the weights $\boldsymbol{\gamma}$. In particular, as $\lambda\to 1/\alpha$ and $\beta\to 0$, the convergence rate gets arbitrarily close to $\mathcal{O}(N^{-\alpha})$.
@@
Here, we consider the worst-case root mean squared $L_{2}$ error by incorporating a random shift. It is natural to ask whether similar optimal convergence rates and good tractability results, the latter of which shall be shown in the next subsection, can be achieved for the deterministic $L_{2}$ error without such a shift. As shown in the proof of Theorem 4.6, for an individual function $f\in H_{d,\alpha,\boldsymbol{\gamma}}$, the squared $L_{2}$ error for the randomized algorithm satisfies
$\int_{[0,1]^{d}}\|f-A_{N,\boldsymbol{\Delta}}^{\mathrm{mult}}(f)\|_{L_{2}}^{2}\,\mathrm{d}\boldsymbol{\Delta}\leq\sum_{\boldsymbol{k}\notin\mathcal{A}_{d}}|\widehat{f}(\boldsymbol{k})|^{2}+\sum_{\boldsymbol{k}\in\mathcal{A}_{d}}\sum_{t=1}^{L}U_{t}(\boldsymbol{k})\sum_{\boldsymbol{\ell}\in P_{n_{t},\boldsymbol{g}_{t}}^{\perp}\setminus\{\boldsymbol{0}\}}|\widehat{f}(\boldsymbol{k}+\boldsymbol{\ell})|^{2},$
@@ 4.3 Tractability
Finally, we discuss the tractability of the multivariate approximation problem in the weighted Korobov space $H_{d,\alpha,\boldsymbol{\gamma}}$. Since Corollaries 4.4 and 4.7 both rely on Proposition A.1, and respectively on Theorems 4.2 and 4.6, the following argument assumes that $M\geq 1$ and $|\mathcal{A}_{d}|\geq 2$. Note that $M$ and $|\mathcal{A}_{d}|$ are closely related through (2). In fact, we require a stronger condition on $M$, i.e., we choose $M\geq\max(1,\min_{u}\gamma_{u}^{-1})$, which ensures $|\mathcal{A}_{d}|\geq 2$ since the corresponding $\mathcal{A}_{d}$ contains $\boldsymbol{0}$ and at least one non-zero vector. Furthermore, we recall that the error bounds in Corollaries 4.4 and 4.7 hold under the assumption that
@@
$\geq 2\max_{j=1,\ldots,d}\max_{j\in u\subseteq[1{:}d]}\lfloor(\gamma_{u}M)^{1/\alpha}\rfloor=N_{\mathcal{A}_{d}},$
@@
Next, to show when the assumption $\eta\geq 4L_{\max}\log L_{\max}$ holds, let us consider the case $\delta=1/2$. Recall that $|\mathcal{A}_{d}|\geq 2$. From the definition of $L_{\max}$, by applying the inequality $\log(x+1)\leq x^{\beta}/\beta$, which holds for any $x\geq 0$ and $\beta\in(0,1)$ (here we take $\beta=1/3$), we have
@@
$\leq 36\sqrt[3]{3}\left(\frac{c}{c-1}\right)^{8/3}\left(1+\frac{1}{\log 2}\right)^{4/3}\left(|\mathcal{A}_{d}|-1\right)^{4/9}.$
@@
$c\geq 36\sqrt[3]{3}\left(\frac{c}{c-1}\right)^{8/3}\left(1+\frac{1}{\log 2}\right)^{4/3}.$
Numerically, this inequality is satisfied for $c\geq 173.46$. It is important to note that the assumptions required for the error bounds in Corollaries 4.4 and 4.7 can be satisfied with $\delta=1/2$ and a constant $c\geq 173.46$ independently of the dimension $d$, smoothness $\alpha$, and the weight parameters $\boldsymbol{\gamma}$.
We now state the conditions for strong polynomial tractability. Following the standard framework in information-based complexity [20, 21, 22], let $N(\varepsilon,d)$ denote the information complexity, which is defined as the minimal number of function evaluations required to reduce the worst-case error to at most $\varepsilon\in(0,1)$ for a $d$-dimensional problem. A multivariate problem is said to be strongly polynomially tractable if there exist absolute constants $C>0$ and $p\geq 0$ such that
$N(\varepsilon,d)\leq C\varepsilon^{-p}\quad\text{for all $\varepsilon\in(0,1)$ and $d\in\mathbb{N}$}.$
In other words, the information complexity is bounded by a polynomial in $\varepsilon^{-1}$ that is completely independent of the dimension $d$. Since we have provided a constructive algorithm that achieves a dimension-independent error bound, the strong polynomial tractability of our problem directly follows. The next result is thus an immediate consequence of Corollaries 4.4 and 4.7.
@@
When the weights $\boldsymbol{\gamma}$ have some specific structure, in particular, when they are either product weights or POD weights, the condition given in the above theorem can be simplified as follows. We refer to the recent work [25] for more refined analyses for the POD and SPOD weights cases.
@@
This completes the proof of the first assertion.
@@
$\sum_{\begin{subarray}{c}u\subset\mathbb{N}\\ |u|=k\end{subarray}}\prod_{j\in u}x_{j}\leq\frac{1}{k!}\left(\sum_{j=1}^{\infty}x_{j}\right)^{k},$
@@
In [8, Theorem 15.11], it is shown that, for product weights, strong polynomial tractability holds if there exists $\tau>1$ such that
@@
This discrepancy stems from the differences in the choice of the index set $\mathcal{A}_{d}$ and the corresponding proof techniques. Our index set $\mathcal{A}_{d}$ is $\alpha$-dependent as in (2), whereas the index set in [8, Theorem 15.11] is $\alpha$-independent. As mentioned in Remark 4.3, the square roots of the weight parameters, $\boldsymbol{\gamma}^{1/2}$, are considered in constructing their index set, which plays an important role in deriving the above $\alpha$-independent summability condition.